OpenStudy (anonymous):

A point charge \(\color{blue}{Q}\) is at the Origin of a Spherical Co-ordinate System. Find the \(\color{blue}{\textsf{Flux}}\) which crosses the portion of a Spherical Shell described by \(\color{green}{a \le \theta \le \beta}\) ??

OpenStudy (perl):

this is calculus 3, do you have the formulas to use?

OpenStudy (anonymous):

Using Gauss's Law, \[\large Net \; \; Flux \; \; crossing \; ( \psi_{net}) = Q_{enclosed}\]

OpenStudy (anonymous):

@aum ..

OpenStudy (aum):

Assume a sphere of radius 'r' enclosing the point charge Q at its center. Net flux crossing the sphere \(\large \psi = Q.\) This flux passes through the entire surface area of the sphere, which is, \(\large 4\pi r^2. \) The flux passing through a portion of a sphere with surface area A will be: \(\large \frac{A}{4\pi r^2} * Q. \) We need to find A. What is the surface area A of a portion of a sphere of radius 'r' bounded by \(\alpha \le \theta \le \beta\) and \(0 \le \phi \le 2\pi\)?

OpenStudy (aum):